# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v3

@@ II-C Signal Recovery Scheme
II-D Proposed Recovery Scheme vs. Other DFT-based Techniques
@@
Due to the DFT computation, our proposed recovery scheme can be classified as a DFT-based recovery technique. Other algorithms under this category are the Fourier-Prony method [10], USLSE [12], $B^{2}R^{2}$ [11], and LASSO-$B^{2}R^{2}$ [13]. The advantage of our proposed recovery scheme is that it does not restrict the signal to decay over time (such as in $B^{2}R^{2}$ and LASSO-$B^{2}R^{2}$) nor does it require the signal to be periodic (such as in Fourier-Prony method). The window function ensures that sharp discontinuities are avoided at the periodic boundaries when the length-$N$ sampled signal is repeated. Consequently, spectral leakage is minimized. The vanishing signal amplitude requirement in $B^{2}R^{2}$ and LASSO-$B^{2}R^{2}$ may require long observation time to estimate the modulo residue samples. The computational complexities of these algorithms grow fast with the observation length. In contrast, the proposed algorithm follows the principle of STFT and the observation length of a segment can be controlled. That is, we can either choose to unfold few overlapping segments with large observation length or many overlapping segments with small observation length. In addition, the sliding DFT approach enables processing of modulo samples in short frames. This is suitable for real-time applications with low latency requirement.
@@
Our objective is to derive MSE performance guarantees for this recovery scheme and demonstrate the advantage of modulo ADCs which uses the proposed recovery scheme over conventional ADCs. In the next section, the main results of our paper are presented.
@@ III-A Measuring Recovery Performance
Our primary measure of recovery performance of the unfolding scheme is the MSE between the (sampled) input signal $f[n]$ and the $\hat{f}[n]$, given by
@@
where $N_{0}$ is the signal length. The second line comes from partitioning the original signal into II length-NN segments. The third line follows from the expression of the unfolded signal in equation (II-C) and from doing some algebraic manipulation. The fourth line comes from the fact that the quantization noise has zero mean. Thus, the second term in the third line can be removed.
From equation (III-A), it can be seen that the MSE expression comes from two sources: (1) the error due to the modulo residue estimation error and (2) the in-band quantization noise. Our objective in this section is to derive the exact MSE expression as a function of $\mathrm{OF}$ and $b$ when our proposed recovery procedure is used.
@@ III-B Statistical Model for Quantization Noise
Since MSE involves the mean square of the quantization noise, it is crucial to understand the statistical properties of the quantization noise. The statistical nature of the quantization noise comes from the dithered quantization framework established in [52]. This framework is appropriate for the modulo ADC since the folding operation ensures that the signal does not overload the quantizer. The ‘no overloading’ property of the modulo ADC, together with the triangle dither $d[n]$, guarantees that the quantization noise is a white process and its power does not depend on the input, i.e. $\mathbb{E}\left\{|\epsilon[n]|^{2}\big{|}f_{\lambda^{\prime}}[n]\right\}=\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}$ (see [52, Theorem 2]). For triangle dither $d[n]$ described in Section II, the induced quantization noise has power $\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}=\frac{1}{4}\left(\frac{2\lambda}{2^{b}}\right)^{2}=\frac{\lambda^{2}}{2^{2b}}$.
@@ III-C Incorporating Spectral Leakage
Since spectral leakage cannot be completely neglected for small $N$, we explicitly account for the effect of non-negligble spectral leakage in the proposed recovery algorithm. To this end, we define the spectral leakage bin count as $K_{\mathrm{SL}}=2\big{\lceil}\frac{\delta_{\mathrm{SL}}\cdot N}{\pi}\big{\rceil}$, which corresponds to the number of discrete frequency bins within the spectral leakage intervals $(\rho\pi,\rho\pi+\delta_{\mathrm{SL}})\cup(2\pi-\rho\pi-\delta_{\mathrm{SL}},2\pi-\rho\pi)$. While $K_{\mathrm{SL}}$ and $\delta_{\mathrm{SL}}$ are generally difficult to compute exactly in practice due to their dependence on the precise shape of the leakage spectrum, we highlight their role in the MSE performance guarantee to reveal how spectral leakage affects recovery accuracy. Finally, we note that the MSE performance guarantee to be derived does not account for the spectral leakage inside $\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}}]$. Nonetheless, our numerical results in Section IV demonstrate that our theoretical MSE predictions still coincide with the simulated MSE despite neglecting some spectral leakage.
@@ III-D Main Result
Before we state the main result of this paper, a proposition about the matrix $\mathbf{V}_{\mathcal{S}_{i}}$ is first presented. This will play a key role in the derivation of the main result.
@@
With $\mathbf{V}_{\mathcal{S}_{i}}$ being full column rank, $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}}$ is an $|\mathcal{S}_{i}|\times|\mathcal{S}_{i}|$ identity matrix. Under a ‘no quantization noise setting’, $\|\underline{\tilde{\mathbf{z}}_{\mathcal{S}_{i}}}-\underline{\mathbf{z}_{\mathcal{S}_{i}}}\|_{\infty}=\|\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}}\underline{{\mathbf{z}}_{\mathcal{S}_{i}}}-\underline{\mathbf{z}_{\mathcal{S}_{i}}}\|_{\infty}=0$. Thus, the full column rank property is crucial for the recovery of $z^{(i)}[n]$.
One key observation to note in Proposition 1 is that the oversampling factor $\mathrm{OF}$ approaches to unity if $|\mathcal{S}_{i}|$ grows strictly slower with $N$. This is true for finite-energy bandlimited functions due to their time-domain decay property (see [13, 11]). That is, $\exists n_{0}$ such that $f[n]<\lambda^{\prime}$ for all $|n|>n_{0}$. Consequently, $\exists N_{0}$ such that $|\mathcal{S}_{i}|$ does not change $\forall N>N_{0}$. Thus, the sampling rate can be made closer to the Nyquist rate for finite-energy bandlimited signals at the expense of longer observation window. Similar trend is observed in the recovery guarantees established in [15]. That is, for signals that are compact $\lambda$-exceedance with parameter $\rho_{e}$, smaller $\mathrm{OF}$ can be used by increasing the number of observations. This $\lambda$-exceedance property is closely related to the time-domain decay property in [13, 11].
The drawback of Proposition 1 is that it depends on the bandlimited input signal $f(t)$. The following proposition provides a sufficient condition for $\mathrm{OF}$ that depends only on the maximum amplitude $\|f(t)\|_{\infty}$ and not on the whole $f(t)$.
@@
Note that Proposition 2 is a sufficient but not a necessary condition. The expression we used to bound $\max_{i}|\mathcal{S}_{i}|$ from above is identical to the bound of the folding instances derived in [13, Equation 8]. However, their proof assumed that $f(t)$ is a bandlimited signal with negligible truncation error. We show in Appendix B that this bound is also valid for the setup considered in this paper. However, this bound can be loose in some cases as demonstrated in [13, Table 1].
For $N$ approaching infinity, equation (13) simplifies to $\mathrm{OF}>3$, which was derived in [14]. The theoretical MSE guarantee established in [14] did not take into account the spectral leakage. As demonstrated in Proposition 2, the spectral leakage bin count $K_{\mathrm{SL}}$ increases the required $\mathrm{OF}$ to satisfy condition (13). Nonetheless, its effect on the OF sufficient condition diminishes fast as $N$ grows.
We can now state the main theoretical result of this paper. The following theorem establishes the MSE of the proposed recovery method when the parameters $\mathrm{OF}$ and $b$ are above certain values.
@@
The intuition behind the result is that the range of $\mathrm{OF}$ and $b$ in Theorem 1 guarantees that the bounded noise added to the modulo residue pre-estimate $\tilde{z}^{(i)}[n]$ is within $(-\lambda^{\prime},+\lambda^{\prime})$. Since $\tilde{z}^{(i)}[n]$ are integer multiples of $2\lambda^{\prime}$, the rounding operation maps all modulo residue pre-estimates $\tilde{z}^{(i)}[n]$ to the correct modulo residue $z^{(i)}[n]$. Consequently, only the quantization noise power term in equation (III-A) appears in the MSE expression.
The derivation of [14, Theorem 1] did not take into account the contribution of the $\epsilon^{(i)}[n]$ for $n\in\mathcal{S}_{i}^{c}$, resulting in $b>3$ condition. In our Theorem 1, the second term in equation (16) accounts for the impact of $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ on the recovery performance. Although we are interested in the modulo residue samples located at $\mathcal{S}_{i}$, $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ still contributes to $\hat{z}^{(i)}[n]$ through the computation of $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$. In Section IV, we investigate how the second term of (16) is affected by parameters such as segment length, cardinality of $\mathcal{S}_{i}$, and oversampling factor.
It is also important to point out the detrimental impact of spectral leakage to the MSE performance. Neglecting spectral leakage, i.e. $\delta_{\mathrm{SL}}=0$, equation (17) simplifies to
@@
This is similar to the asymptotic growth rate established in [14] which neglected spectral leakage. When $\delta_{\mathrm{SL}}>0$, the asymptotic growth rate of the MSE becomes $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ only. $\delta_{\mathrm{SL}}$ can be reduced if larger window length is used. However, using a larger window length increases the computational complexity of the sliding window DFT (see Section III-G).
@@ III-E Comparison with Recovery Guarantees of Existing Methods
We now compare the derived theoretical results with other performance guarantees derived for modulo sampling. One popular recovery procedure is the higher-order differences (HoD) approach developed in [7]. Performance guarantee for this algorithm under a bounded noise setting (e.g., quantization noise) has been analyzed in [7, Theorem 6]. More precisely, they showed that for finite-energy bandlimited signals, the noisy unfolded samples can be recovered from noisy modulo samples up to an unknown additive constant, i.e.,
$\hat{f}_{\mathrm{HoD}}[n]=f[n]+\epsilon[n]+2\lambda^{\prime}p,$
@@
where $p\in\mathbb{Z}$ is not known, if the sampling rate is at least $2^{\alpha}\pi e\times f_{\mathrm{Nyq}}$. Here, $\alpha\in\mathbb{N}$ is a parameter of the bounded noise and depends on both the maximum amplitude of the noise and the modulo threshold. In contrast, Theorem 1 only requires $\mathrm{OF}>3$ under negligible spectral leakage. We also note that $\mathrm{OF}>3$ closely resembles the OF requirement derived in [53] to uniquely identify periodic bandlimited signals under the modulo-DFT sensing model. However, their model first applies DFT on the sampled signal prior to modulo operation. Moreover, the $\mathrm{OF}>3$ result is equivalent to Itoh’s condition [54] and is reminiscent of the spatial oversampling requirement established for DoA estimation using MIMO arrays with modulo nonlinearities [33].
@@
A prediction-based approach [16] showed that $\mathrm{OF}$ can be made arbitrarily close to unity in exchange for a significant increase in prediction filter length. However, it was pointed out in [16, Section III] that their proposed approach “collapses in the presence of quantization noise”. In contrast, our algorithm can work with finite quantization bits. Moreover, our proposed algorithm can work at a sampling rate arbitrarily close to the Nyquist rate for finite-energy bandlimited signals and sufficiently large window length $N$ by virtue of Proposition 1. We also note that the $\mathrm{OF}$ requirement established in [10] for periodic bandlimited signals is
$\mathrm{OF}\geq\frac{N}{N-2|\mathcal{S}|-2},$
@@
where $|\mathcal{S}|$ is the number of folding instances. Evidently, when $K_{\mathrm{SL}}=0$, the sufficient condition on $\mathrm{OF}$ established in Proposition 1 is better than the $\mathrm{OF}$ requirement in [10]. The advantage of Proposition 1 might be attributed to the availability of the 1-bit folding information in our setup.
@@
Finally, there exists a line of work [28, 29, 31] which uses time-domain thresholding to unfold the output produced by generalized modulo sampling models. Unlike the DFT-based recovery methods, thresholding is done in the time domain so spectral leakage is not an issue. Moreover, thresholding-based techniques have established recovery guarantees [28, 29]. However, the sequential estimation of the folding parameters requires that the folding times are well-separated in the time domain. Such separation cannot be guaranteed under ideal modulo sampling with zero hysteresis as demonstrated in [29, Lemma 1]. Moreover, the MSE bound established in [29, Proposition 1] is inversely proportional to the hysteresis $h$. Under ideal modulo nonlinearities, this $h$ goes to zero and so the MSE under the thresholding-based recovery becomes unbounded.

### v4

@@ II-D Design Considerations for Window Function
II-E Proposed Recovery Scheme vs. Other Fourier-based Reconstruction Techniques
@@
Because of the STFT computation, our proposed recovery scheme can be classified as a Fourier-based recovery technique. Other algorithms under this category are the Fourier-Prony method [6], USLSE [53], $B^{2}R^{2}$ [1], and LASSO-$B^{2}R^{2}$ [45]. The advantage of our proposed recovery scheme is that it does not restrict the signal to decay over time (such as in $B^{2}R^{2}$ and LASSO-$B^{2}R^{2}$) nor does it require the signal to be periodic (such as in the Fourier-Prony method). The window function ensures that sharp discontinuities are avoided at the periodic boundaries when the length-$N$ sampled signal is repeated. Consequently, spectral leakage is minimized. The vanishing signal amplitude requirement in $B^{2}R^{2}$ and LASSO-$B^{2}R^{2}$ may require a long observation time to estimate modulo residue samples. The computational complexities of these algorithms increase rapidly with the observation length. In contrast, the proposed algorithm follows the principle of STFT and the observation length of a segment can be controlled. As such, we can either choose to unfold few overlapping segments with large observation length or many overlapping segments with small observation length. In addition, the STFT-based approach enables processing of modulo samples in short frames. This is suitable for real-time applications with low-latency requirements.
@@
Our objective is to derive the MSE performance guarantees and computational complexity for this recovery scheme and demonstrate the advantage of modulo ADCs that use the proposed recovery scheme over conventional ADCs. The main results of our paper are presented in the next section.
@@ III Performance Guarantees using the Proposed Recovery Procedure
In this section, we establish theoretical performance guarantees for the proposed STFT-based signal recovery method. We begin by introducing the primary performance metric, the mean squared error (MSE), and justify the statistical model for the dither signal and resulting quantization noise. Next, we define the parameters that capture the effects of (non-negligible) spectral leakage in the DFT computation. Building on these foundations, we present sufficient conditions for the oversampling factor and quantization resolution to ensure accurate unfolding of modulo residues. We also compare the performance of the proposed STFT-based recovery method with that of conventional ADCs and existing modulo ADC recovery techniques. The section concludes with a discussion of the proposed algorithm’s computational complexity.
@@ III-A Measuring Recovery Performance
Our primary measure of the recovery performance of the unfolding scheme is the MSE between the (sampled) input signals $f[n]$ and the $\hat{f}[n]$, given by
@@
where $N_{0}$ denotes the signal length. The second line was obtained from partitioning the original signal into II length-NN segments. The third line follows from the expression of the unfolded signal in equation (II-C) and from performing some algebraic manipulation. The fourth line comes from the fact that the quantization noise has zero mean. Thus, the second term on the third line can be removed.
From equation (III-A), it can be seen that the MSE expression comes from two sources: (1) the error due to the modulo residue estimation error and (2) the in-band quantization noise. Our objective in this section is to derive the exact MSE expression as a function of $\mathrm{OF}$ and $b$ when the proposed recovery procedure is used.
@@ III-B Statistical Model for Quantization Noise
Since MSE involves the mean square of the quantization noise, it is crucial to understand the statistical properties of the quantization noise. The statistical nature of the quantization noise originates from the dithered quantization framework established in [25]. This framework is appropriate for modulo ADC because the folding operation ensures that the signal does not overload the quantizer. The ‘no overloading’ property of the modulo ADC, together with the dither signal $d[n]$, guarantees that the quantization noise is a white process.
In our problem setup, we use a triangular dither rather than a uniform dither because of its favorable second-order statistical properties. In particular, the power of the quantization noise induced by the triangular dither is independent of the input, i.e.,
$\mathbb{E}\left\{|\epsilon[n]|^{2}\big|f_{\lambda^{\prime}}[n]\right\}=\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}$ (14)
as established in [25, Theorem 2]. This input-independence of the second-order statistics of quantization noise does not generally hold for a single uniform dither [25]. For the triangular dither $d[n]$ described in Section II, the resulting quantization noise power is
$\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}=\frac{1}{4}\left(\frac{2\lambda}{2^{b}}\right)^{2}=\frac{\lambda^{2}}{2^{2b}}$ (15)
@@ III-C Incorporating Spectral Leakage
Since spectral leakage cannot be completely neglected for small $N$, we explicitly accounted for the effect of non-negligible spectral leakage in the proposed recovery algorithm. Thus, we define the spectral leakage bin count as
$K_{\mathrm{SL}}=2\bigg\lceil\frac{\delta_{\mathrm{SL}}\cdot N}{\pi}\bigg\rceil,$ (16)
which corresponds to the number of discrete frequency bins within the spectral leakage interval $(\rho\pi,\rho\pi+\delta_{\mathrm{SL}})\cup(2\pi-\rho\pi-\delta_{\mathrm{SL}},2\pi-\rho\pi)$. The frequency-domain plots in Figure 4 provides a visual illustration of the intuition behind the spectral leakage bin count. Although $K_{\mathrm{SL}}$ and $\delta_{\mathrm{SL}}$ are generally difficult to compute exactly in practice due to their dependence on the precise shape of the leakage spectrum, we highlight their role in the MSE performance guarantee to reveal how spectral leakage affects recovery accuracy. Finally, we note that the MSE performance guarantee to be derived does not account for the spectral leakage inside $\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}}]$. Nonetheless, our numerical results in Section IV demonstrate that our theoretical MSE predictions still coincide with the simulated MSE despite neglecting some spectral leakage.
@@ III-D Main Result
Before we state the main result of this paper, a proposition about the matrix $\mathbf{V}_{\mathcal{S}_{i}}$ is presented. This plays a key role in deriving the main result.
@@
With $\mathbf{V}_{\mathcal{S}_{i}}$ being full column rank, $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}}$ is a $|\mathcal{S}_{i}|\times|\mathcal{S}_{i}|$ identity matrix. Under a ‘no quantization noise setting’, $\|\underline{\tilde{\mathbf{z}}_{\mathcal{S}_{i}}}-\underline{\mathbf{z}_{\mathcal{S}_{i}}}\|_{\infty}=\|\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}}\underline{{\mathbf{z}}_{\mathcal{S}_{i}}}-\underline{\mathbf{z}_{\mathcal{S}_{i}}}\|_{\infty}=0$. Thus, the full column rank property is crucial for the recovery of $z^{(i)}[n]$.
One key observation in Proposition 1 is that the oversampling factor $\mathrm{OF}$ approaches unity if $|\mathcal{S}_{i}|$ grows strictly slower with $N$. This is true for finite-energy bandlimited functions due to their time-domain decay property (see [45, 1]). That is, $\exists n_{0}$ such that $f[n]<\lambda^{\prime}$ for all $|n|>n_{0}$. Consequently, $\exists N_{0}$ such that $|\mathcal{S}_{i}|$ does not change $\forall N>N_{0}$. Thus, the sampling rate can be made closer to the Nyquist rate for finite-energy bandlimited signals at the expense of a longer observation window. Similar trend was observed in the recovery guarantees established in [2]. That is, for signals that are compact $\lambda$-exceedance with parameter $\rho_{e}$, smaller $\mathrm{OF}$ can be used by increasing the number of observations. This $\lambda$-exceedance property is closely related to the time-domain decay property in [45, 1].
One drawback of Proposition 1 is that it depends on the bandlimited input signal $f(t)$. The following proposition provides a sufficient condition for $\mathrm{OF}$ that depends only on the maximum amplitude $\|f(t)\|_{\infty}$ and not on the entire signal $f(t)$.
@@
Note that Proposition 2 is sufficient, but not a necessary condition. The expression used to bound $\max_{i}|\mathcal{S}_{i}|$ from above is identical to the bound of the folding instances derived in [45, Equation 8]. However, their proof assumed that $f(t)$ is a bandlimited signal with negligible truncation error. We show in Appendix B that this bound on $\max_{i}|\mathcal{S}_{i}|$ is also valid for the setup considered in this paper. However, this bound can be loose in some cases as demonstrated in [45, Table 1].
For $N$ approaching infinity, equation (18) simplifies to $\mathrm{OF}>3$, which was derived in [4]. The theoretical MSE guarantee established in [4] does not take into account the spectral leakage. As demonstrated in Proposition 2, the spectral leakage bin count $K_{\mathrm{SL}}$ increases the $\mathrm{OF}$ required to satisfy condition (18). Nonetheless, its effect on the OF sufficient condition diminishes fast as $N$ grows.
We now state the main theoretical result of this paper. The following theorem establishes the MSE of the proposed recovery method when parameters $\mathrm{OF}$ and $b$ are above certain values.
@@
The intuition behind this result is that the sufficient conditions for $\mathrm{OF}$ and $b$ in Theorem 1 guarantee that the bounded noise added to the modulo residue pre-estimate $\tilde{z}^{(i)}[n]$ is within $(-\lambda^{\prime},+\lambda^{\prime})$. Since $\tilde{z}^{(i)}[n]$ are integer multiples of $2\lambda^{\prime}$, the rounding operation maps all modulo residue pre-estimates $\tilde{z}^{(i)}[n]$ onto the correct modulo residue $z^{(i)}[n]$. Consequently, only the quantization noise power term in equation (III-A) appears in the MSE.
The derivation of [4, Theorem 1] did not consider the contribution of the $\epsilon^{(i)}[n]$ for $n\in\mathcal{S}_{i}^{c}$, resulting in $b>3$ condition. In Theorem 1, the second term in equation (21) accounts for the impact of $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ on recovery performance. Although we are interested in the modulo residue samples located at $\mathcal{S}_{i}$, $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ still contributes to $\hat{z}^{(i)}[n]$ through the computation of $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$. In Section IV, we investigate how the second term of (21) is affected by parameters such as segment length, cardinality of $\mathcal{S}_{i}$, and oversampling factor.
It is also important to point out the detrimental impact of spectral leakage on MSE performance. Neglecting the spectral leakage, that is, $\delta_{\mathrm{SL}}=0$, equation (22) simplifies to
@@
This is similar to the asymptotic growth rate established in [4] which neglected spectral leakage. When $\delta_{\mathrm{SL}}>0$, the asymptotic growth rate of MSE becomes $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ only. $\delta_{\mathrm{SL}}$ can be reduced by using a larger window length. However, using a larger window length increases the computational complexity of sliding window DFT (see Section III-G).
@@ III-E Comparison with Recovery Guarantees of Existing Methods
We now compare the derived theoretical results with other performance guarantees derived for modulo sampling. A popular recovery procedure is the higher-order differences (HoD) approach developed in [9]. Performance guarantee for this algorithm under a bounded noise setting (e.g., quantization noise) was analyzed in [9, Theorem 6]. More precisely, they showed that for finite-energy bandlimited signals, noisy unfolded samples can be recovered from noisy modulo samples up to an unknown additive constant, that is,
$\hat{f}_{\mathrm{HoD}}[n]=f[n]+\epsilon[n]+2\lambda^{\prime}p,$ (25)
@@
where $p\in\mathbb{Z}$ is unknown, if the sampling rate is at least $2^{\alpha}\pi e\times f_{\mathrm{Nyq}}$. Here, $\alpha\in\mathbb{N}$ is a parameter of the bounded noise and depends on both the maximum amplitude of the noise and modulo threshold. In contrast, Theorem 1 only requires $\mathrm{OF}>3$ under negligible spectral leakage. We also note that $\mathrm{OF}>3$ closely resembles the OF requirement derived in [55] to uniquely identify periodic bandlimited signals under the modulo-DFT sensing model. However, their model first applies DFT to the sampled signal prior to modulo operation. Moreover, the $\mathrm{OF}>3$ result is equivalent to Itoh’s condition [28] and is reminiscent of the spatial oversampling requirement established for DoA estimation using MIMO arrays with modulo nonlinearities [13].
@@
A prediction-based approach [43] showed that $\mathrm{OF}$ can be made arbitrarily close to unity in exchange for a significant increase in the prediction filter length. However, it was pointed out in [43, Section III] that their proposed approach “collapses in the presence of quantization noise”. In contrast, our algorithm can operate with finite quantization bits. Moreover, our proposed algorithm works at sampling rates arbitrarily close to the Nyquist rate for finite-energy bandlimited signals and a sufficiently large window length $N$ by virtue of Proposition 1. We also note that the $\mathrm{OF}$ requirement established in [6] for periodic bandlimited signals is
$\mathrm{OF}\geq\frac{N}{N-2|\mathcal{S}|-2},$ (26)
@@
where $|\mathcal{S}|$ denotes the number of folding instances. Evidently, when $K_{\mathrm{SL}}=0$, the sufficient condition for the $\mathrm{OF}$ established in Proposition 1 is better than the $\mathrm{OF}$ requirement in [6]. The advantage of Proposition 1 might be attributed to the availability of 1-bit folding information in our setup.
@@
Finally, there exists a line of work [19, 16, 29] that uses time-domain thresholding to unfold the output produced by generalized modulo sampling models. Unlike DFT-based recovery methods, thresholding is performed in the time domain; therefore, spectral leakage is not an issue. Moreover, thresholding-based techniques have established recovery guarantees [19, 16]. However, the sequential estimation of the folding parameters requires that the folding times be well-separated in the time domain. Such a separation cannot be guaranteed under ideal modulo sampling with zero hysteresis as demonstrated in [16, Lemma 1]. Moreover, the MSE bound established in [16, Proposition 1] is inversely proportional to the hysteresis $h$. Under ideal modulo nonlinearities, this $h$ goes to zero, and thus the MSE under the thresholding-based recovery becomes unbounded.